OpenStudy (anonymous):

Find the equation of the line containing the point (-2,3) and the slope of 0

sam (.sam.):

y=3

OpenStudy (saifoo.khan):

y = 3

OpenStudy (anonymous):

How?

sam (.sam.):

\[y-y _{1}=m(x-x _{1})\] \[x=-2,y=3,m=0\] \[y-3=0\]\[y=3\]

sam (.sam.):

Correction \[x _{1}=−2,y _{1}=3,m=0\]